### v2

# Integration vs segregation: evidence from measuring and comparing interdisciplinarity in funded and unfunded research using a network approach

Abstract
Interdisciplinary research fuels innovation. However, research investment has been found to have become more conservative and highly interdisciplinary ideas tend to be punished during research evaluation. How does this affect the interdisciplinarity of research output? To answer this question, our study presents a novel network approach by comparing the evolution of interdisciplinarity in funded research against unfunded research. Considering the field of infectious disease research, the analysis of temporal networks reveals (i) the growing compartmentalisation of funded research, i.e., it becomes more conservative and focuses on the groups of knowledge with readily established connections, and (ii) the growth in global integration in unfunded research, i.e., it tends to be more widely exploratory and links distant knowledge. In addition, we find that in general funded research is less interdisciplinary than unfunded research but this effect decreases through time. Moreover, we observe that interdisciplinary research on prominent infectious diseases like HIV tends to have strong bridging effects facilitating global knowledge integration in the network. In terms of the coronavirus research, despite the surge in publications since COVID-19, we find its systematic impact on knowledge integration remains relatively low. Overall, this research provides detailed insights into the impact of conservatism in funding interdisciplinary research, allows for the identification of under/over-funded interdisciplinary research areas, and offers valuable implications for policymakers aiming to foster interdisciplinary research.
keywords:
Interdisciplinarity , Knowledge Integration , Funding , Integration , Segregation , Temporal Network
journal: Journal of Informetrics
[…]
[label1]organization=School of Electronic and Computer Science, University of Southampton, city=Southampton, postcode=SO17 1BJ, country=UK
[…]
[label2]organization=Faculty of Medicine, University of Southampton, city=Southampton, postcode=SO17 1BJ, country=UK
{highlights}
We propose a novel framework to extract and compare funded and unfunded interdisciplinary research in any field of interest and identify potentially under/over-funded interdisciplinary research areas.
We show that conservatism in research funding leads to a compartmentalised structure of funded interdisciplinary research, i.e., established clusters of knowledge get reinforced. On the other hand, unfunded research is generally more exploratory and links more distant fields of study.
We examine how important real-world events influence interdisciplinary research, i.e., we study how interdisciplinary research efforts react to the 2002-2004 SARS outbreak, the 2015-2016 zika outbreak and the 2019-2022 COVID-19 pandemic.
1 Introduction
Interdisciplinary research (IDR), a process of knowledge integration (Rafols and Meyer, 2010), has been seen as a source of creativity and innovativeness (Rousseau et al., 2019). Various policy and funding initiatives have been developed to facilitate IDR (Wang and Shapira, 2015). However, it has been discovered that highly interdisciplinary (Bromham et al., 2016) and highly novel (Boudreau et al., 2016) ideas that integrate knowledge in unprecedented ways (Uzzi et al., 2013) tend to be penalised during grant proposal evaluation. Given such a lack of appreciation towards highly interdisciplinary ideas at the stage of funding application, does it carry further to research output generated by funding? It is an important question to ask as proposals represent the intention of a scientific idea while publications represent the realisation (Packalen and Bhattacharya, 2020), i.e., when the intention of knowledge integration is discouraged, what happens to the realisation of knowledge integration?
Funding’s influences have been explored in a wide range of aspects, including on research methods and goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), scientific productivity (Hottenrott and Lawson, 2017), and scholarly, technological and social impact of the associated research output (Heyard and Hottenrott, 2021; Yang, 2024). Of those studies examining various aspects related to research output produced by funding (Jacob and Lefgren, 2011; Wahls, 2019; Packalen and Bhattacharya, 2020; Arora and Gambardella, 2006; Heyard and Hottenrott, 2021; Arora et al., 2000; Benavente et al., 2012; Yang, 2024), limited attention has been paid to examine its interdisciplinarity (ID). In addition, studying funded research output beyond individual funder level at scale remains a gap (Yang, 2024).
In this paper, we address the above gaps. By identifying all funded and unfunded publications in the field of infectious disease from 1995 to 2022, we propose a novel temporal network approach to characterise and compare the ID of funded and unfunded research. We aim to answer the following questions:
  1. 1.
    Are funded research less interdisciplinary in general compared to unfunded research? What are the most under/over-funded interdisciplinary research areas?
  2. 2.
    What are the main trends and tendencies in the time-dependence of ID in funded and unfunded research?
  3. 3.
    Can we identify different temporal regimes in the evolution of ID in funded and unfunded research?
  4. 4.
    What roles does research into prominent infectious diseases like HIV and coronavirus play in terms of interdisciplinary knowledge generation?
  5. 5.
    What is the effect of important events like the 2002-2004 SARS outbreak or the COVID-19 pandemic on the ID of funded and unfunded infectious disease research?
The paper is structured as follows: Section 2 provides a brief review of research on ID and funded research, Section 3 introduces the data source and computational implementation of IDR measures, Section 4 presents the main results and Section 5 summarises the results and discusses real-world implications, limitations, and future work.
2 Literature Review
We briefly review research on IDR in Section 2.1 and the central concepts of quantitative IDR measures (diversity, coherence and intermediation) in Section 2.2; then, we will review research on funded research and highlight the gaps in the field in Section 2.3.
2.1 Interdisciplinary Research: Overview
IDR, a process of knowledge integration, generates creativity and innovativeness (National Academy of Sciences et al., 2005; Rafols and Meyer, 2010; Rousseau et al., 2019). Current research on IDR has concluded that IDR could take place cognitively or socially (Glänzel and Debackere, 2022), meaning the integration either happens in a scientist’s mind (Rafols and Meyer, 2010) or in a social process (Abramo et al., 2012, 2017), i.e., formation of a team of researchers with different expertise, experiences, or academic background. Based on these two perspectives, researchers have been exploring a broad range of questions on IDR: from the quantification of the ID of articles (Rafols and Meyer, 2010), journals (Leydesdorff, 2007; Rodríguez, 2017; Leydesdorff et al., 2018), scientists (Porter et al., 2007), institutes (Rafols et al., 2012; Soós and Kampis, 2012), grant proposals (Bromham et al., 2016; Nichols, 2014), and research fields (Leydesdorff and Rafols, 2011), to mapping the global structure of science (Leydesdorff and Rafols, 2009; Rafols et al., 2010), further to the scholarly and social impact of IDR (Okamura, 2019; Hu et al., 2024; Shi et al., 2009; Wang et al., 2015; Yegros-Yegros et al., 2015). For example, Rafols and Meyer (2010) used diversity and coherence as a framework to compare the ID of a set of bionanoscience articles. Leydesdorff and Rafols (2009) and Leydesdorff (2007) analysed the ID of journals based on betweenness centrality and diversity. Porter et al. (2007) proposed the measures of integration, reach and specialisation and used them to measure the ID of $47$ researchers. Soós and Kampis (2012) analysed and compared the disciplinary structure of a sample of Hungarian Research Institutions based on the science overlay maps (Rafols et al., 2010).
IDR may lack institutional appreciation due to mono-disciplinary academic structures (Rousseau et al., 2019) and is poorly rewarded by funders (Bromham et al., 2016) due to disciplinary-based evaluations (Rousseau et al., 2019; Woelert and Millar, 2013). Rafols et al. (2012) compared the ID of Innovation Studies (IS) units with leading Business & Management Schools (BMS) in the UK. Despite IS units being more interdisciplinary, they were disadvantaged in evaluation and obtaining resources due to the bias of the Association of Business Schools’ (ABS) journal rankings favouring mono-disciplinary research. IDR has been found to exhibit higher scholarly impact (Okamura, 2019) but Shi et al. (2009), Wang et al. (2015) and Yegros-Yegros et al. (2015) reported mixed results across different measures (Wang et al., 2015; Yegros-Yegros et al., 2015) and different fields (Shi et al., 2009). IDR has also been found to attract more attention from policy documents (Hu et al., 2024).
2.2 Central Concepts in Quantitative Research on IDR
In IDR studies considerable effort has been used to develop quantitative measures to inform policymakers, research managers, evaluators and sociologists of science (Wagner et al., 2011). Here we briefly revisit relevant concepts in the quantification of IDR, but see Wagner et al. (2011) and Wang and Schneider (2020) for more comprehensive reviews.
There are three central concepts that quantitative IDR measures have been based on: diversity, coherence, and intermediation. Diversity refers to the difference in the bodies of knowledge that are integrated, and it consists of variety, balance, and disparity (Stirling, 2007). The Rao-Stirling (RS) index is one of the most widely-used diversity-based metrics (Wang and Schneider, 2020; Nichols, 2014), and Zhang et al. (2016) proposed the Hill-type measure as an improvement to the low discriminating power of the RS index. In addition, diversity measures adapted from evolutionary biology (Bromham et al., 2016) have also been used to measure ID, e.g., Bromham et al. (2016) used the Phylogenetic Species Evenness (PSE), a measure of the biodiversity of species, to capture ID.
Coherence describes the extent of relatedness of bodies of knowledge (Rafols and Meyer, 2010; Rafols, 2014). Coherence can be defined and used differently (Rafols and Meyer, 2010; Soós and Kampis, 2012; Rafols et al., 2012), and measures of coherence (and also intermediation) are still at an exploratory stage (Rafols et al., 2012). When coherence was first proposed (Rafols and Meyer, 2010), it was used to measure the overall compactness of an article’s knowledge structure. Rafols used a bibliographic coupling network to represent the knowledge structure. The nodes were the article’s references and the links measured connection strengths (based on shared references of references). The coherence of each article is then computed as the mean linkage strength of the network. In contrast, Soós and Kampis (2012) focused on the coherence of research institutions, expressed by the sum of weighted distances of the Web of Science (WOS) Subject Category [^1] Formerly known as the Institute for Scientific Information (ISI) Subject Category. (SC) in an institute’s publication profile. Distances between WOSSCs are based on the global map of science (Leydesdorff and Rafols, 2009), and the weights assigned to the distances are the intensity of interactions between WOSSCs in an institute’s publication profile, i.e., cooccurrences. Soós and Kampis (2012) further suggested multimodality could also reflect coherence, where multimodality is defined as the size distribution of connected components of the network of SCs and quantified by the Shannon-Wiener entropy (Shannon, 1948). Rafols (2014) later proposed tentatively that the concept of coherence of interconnected bodies of knowledge consists of density (number of links), intensity (strength of links) and disparity (degree of difference in two bodies of knowledge that links bridge), but its added value and feasibility remain uncertain (Rousseau et al., 2019).
Intermediation refers to the ability to link distant bodies of knowledge (Leydesdorff, 2007). Betweenness was first proposed by Leydesdorff (2007) as a measure of intermediation. Leydesdorff thinks journals with high betweenness are more interdisciplinary due to their capability to relate otherwise non-interacting journals. The average Local Clustering Coefficient (LCC) was also used to capture the effect of intermediation (Rafols et al., 2012; Soós and Kampis, 2012). The LCC of a body of knowledge reflects to what extent its neighbouring fields are directly connected (or integrated), indicating the tendency of transitivity in the neighbouring field, i.e., to what extent intermediation “results” in integration (Soós and Kampis, 2012). In Rafols et al. (2012), this concept was operationalised slightly differently by assigning weights (the proportion of publications from each body of knowledge) to each LCC. Soós and Kampis (2012) further proposed the network diameter (or the maximal shortest path length) to capture both coherence and intermediation. A low diameter means nodes are quickly reachable from each other, i.e., highly coherent. On the other hand, a high diameter indicates some constituent fields are distant but are still linked through a “mediator”, implying the role of intermediation.
2.3 Funded research and interdisciplinarity
The influence of funding on research has been studied in various aspects (see Thelwall et al. (2023) for further detail), including research methods or goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), researcher, team and institutional level productivity (Heyard and Hottenrott, 2021; Hottenrott and Lawson, 2017), output types (Thelwall et al., 2023), and scholarly, technological and social impacts of outputs (Heyard and Hottenrott, 2021; Yang, 2024).
Studies on funded research output have been mostly based on publications associated with individual funders, e.g., the National Institutes of Health (NIH) (Jacob and Lefgren, 2011; Wahls, 2019; Packalen and Bhattacharya, 2020; Yang, 2024), the National Science Foundation (NSF) (Arora and Gambardella, 2006; Yang, 2024), the Swiss National Science Foundation (Heyard and Hottenrott, 2021), the Italian National Research Council (Arora et al., 2000), the Chilean National Science and Technology Research Fund (Benavente et al., 2012). The recent development of the SciSciNet (Lin et al., 2023) has opened the possibility for larger-scale analyses of research output and outcome generated by NIH and NSF funding (Yang, 2024). Yet, there are still two open questions: (i) limited attention has been paid to thoroughly study the ID of funded research output as the existing works mainly focused on funding proposals (Bromham et al., 2016; Nichols, 2014), and (ii) studying a more complete set of funded research beyond individual funder level remains a gap (Yang, 2024).
We address these gaps by identifying and comparing all funded and unfunded publications in the field of infectious disease research in the WOS. By regarding funded and unfunded research as two dynamic processes and portraying the development of their ID from a temporal network perspective, our work also addresses the frequently stressed limitation that existing analyses of IDR are often static (Rafols and Meyer, 2010; Wagner et al., 2011; Wang and Schneider, 2020; Rousseau et al., 2019). Lastly, categorising papers based on journal classification systems causes inaccuracies (Wagner et al., 2011; Rafols, 2014; Rousseau et al., 2019) and such systems are often too coarse to capture knowledge integration at finer levels (Rafols et al., 2012). Therefore, instead of using journal classification systems like the WOSSC, we conduct topic-level analysis based on a selected set of major infectious diseases.
3 Data and Methods
3.1 Data source and curation
Infectious diseases continue to be a major global health issue with significant impact on human society (Baker et al., 2022). Following one of our co-author’s work Head et al. (2020), we based our study on research into 34 major infectious diseases. [^2] The authors of Head et al. (2020) selected a group of 37 prominent infectious diseases for their funding versus disease burden analysis. Our selection excluded sexually transmitted infections and enteric infections because they are a higher level concept than others. We also added consideration of causal relations: trachoma is caused by chlamydia trachomatis but chlamydia also happens to be in the set, so we excluded trachoma. In addition here, we included coronavirus and diphtheria. The inclusion of coronavirus is due to its recent prominent role in the field of infectious disease research. The inclusion of diphtheria is because two other highly related terms, pertussis and tetanus, have already been included in the group (the DTP combination vaccine prevents against all three diseases and they are thus commonly considered together; the three diseases will be addressed by DTP onwards). This gives us 36 selected infectious diseases, which account for \$70 billion infectious disease research funding (65% of overall total) from G20 countries between 2000 and 2017 (Head et al., 2020), and 94% of the total infectious disease burden [^3] Measured in Disability-Adjusted Life Years (DALYs). One DALY represents the loss of the equivalent of one year of full health. in the 2021 Global Burden of Disease data [^4] Available from https://vizhub.healthdata.org/gbd-results/. (Date last accessed: 2024.9.4).
We selected the Web of Science (WoS) Core Collection database as our data source due to its extensive usage in the scientometrics studies (Birkle et al., 2020). We extracted, for each of the 36 diseases, both funded research and all research, the number of records containing the disease in their title, abstract, or author keywords, i.e., the number of occurrences. We also extracted, for both funded and all research, the number of records containing each possible pair of diseases in their title, abstract, or author keywords, i.e., the number of co-occurrences. To retrieve maximum funded records efficiently, we adopt the method of the right-hand truncation search strategy in both the funding agency (FO) and funding grants (FG) field suggested in Liu et al. (2020), as detailed in Search Query 1. We refer to papers that have FO or FG information in the WoS as funded research (labelled “F”) [^5] All research will be labelled “A” hereafter. and those that don’t as unfunded research (labelled “U”). We searched only the document types of original research article, data paper, and proceeding paper, and we specifically excluded any withdrawn or retracted publications, as detailed in Search Query 2. The timespan considered in this paper is from 1995 to 2022 inclusive [^6] The WoS did not include the author keywords for articles until 1991. The years 1991-1994 were excluded from our analysis as the link profiles of these years were very volatile and did not seem to form coherent clusters with any of the rest of the years like the ones in Figure 1.. See Table Data availability for the search strategy for each disease [^7] Search terms for each disease are firstly picked based on experts’ advice and the entry terms (a list of synonyms) in the Medical Subject Heading (MeSH) database. A further selection of search terms was done after testing the added benefit of the search terms, i.e., search terms with low or no additionally identified papers were dropped. [^8] The search term of herpes in our study contains infections related to Herpes Simplex Viruses (HSV) and Shingles. as well as the total and funded number of records returned.
Search Query 1: FO = (A* OR B* OR C* OR D* OR E* OR F* OR G* OR H* OR I* OR J* OR K* OR L* OR M* OR N* OR O* OR P* OR Q* OR R* OR S* OR T* OR U* OR V* OR W* OR X* OR Y* OR Z* OR 0* OR 1* OR 2* OR 3* OR 4* OR 5* OR 6* OR 7* OR 8* OR 9*) OR FG = (A* OR B* OR C* OR D* OR E* OR F* OR G* OR H* OR I* OR J* OR K* OR L* OR M* OR N* OR O* OR P* OR Q* OR R* OR S* OR T* OR U* OR V* OR W* OR X* OR Y* OR Z* OR 0* OR 1* OR 2* OR 3* OR 4* OR 5* OR 6* OR 7* OR 8* OR 9*)
Search Query 2: DT = ((Article OR Data Paper OR Proceedings Paper) NOT (Retracted Publication OR Withdrawn Publication OR Retraction))
The extracted occurrences and cooccurrences of infectious diseases were recorded in tensors $(C_{F})_{36\times 36\times 28}$ for F and $(C_{A})_{36\times 36\times 28}$ for A. An entry $c_{ij}^{t}$ of $C_{F}$ (or $C_{A}$) gives the cooccurrence count between disease $i$ and $j$ in F (or A) in time slice $t$ [^9] The 28 time slices correspond to years from 1995 to 2022. if $i\neq j$, and the occurrence of $i$ otherwise. The tensor of U $(C_{U})_{36\times 36\times 28}$ is a direct result of element-wise subtraction between $C_{A}$ and $C_{F}$. The data extraction was performed using the Web of Science API Lite [^10] The Web of Science API Lite: support search and data integration using Web of Science data returned as JSON or XML. https://developer.clarivate.com/apis/woslite (Date last accessed: 2024.9.4) in Python. For abbreviations of diseases used in this study, please refer to Table Data availability.
We used one of the most widely-used normalisation measures in scientometrics to normalise the cooccurrence matrix, namely the cosine (Eck and Waltman, 2009) (also known as the Ochiai index (Zhou and Leydesdorff, 2016) or Salton’s index (Adnani et al., 2020)), i.e., $w_{ij}^{t}=\frac{c_{ij}^{t}}{\sqrt{c_{ii}^{t}c_{jj}^{t}}}$. In our setting, the cosine normalises the cooccurrences between $i$ and $j$ at time slice $t$ by the geometric mean of the occurrences of $i$ and $j$ at $t$ when $i\neq j$, and thus measures the correlation between numbers of publications about diseases. When $i=j$, $w_{ij}^{t}=0$. By normalising $C_{A}$, $C_{F}$ and $C_{U}$, we ended up with $W_{A}$, $W_{F}$, $W_{U}$, based on which we then performed the following analysis. Note that we regarded the collection of all WoS publications related to infectious diseases as the system of infectious disease knowledge, where $w_{ij}^{t}$ is the association, or the cognitive proximity, or the current extent of knowledge integration between disease $i$ and $j$ at time $t$.
We also note an important limitation of using the WoS for funding analysis: funded research is under-represented in the database. The WoS has begun to collect funding information since Aug 2008, and the assignment of funding information of publications before this time has been done retrospectively (Liu et al., 2020), so likely a lesser proportion of funded papers have been identified pre-2008 compared to post-2008 as can be seen from Table 1. However, we can see research on infectious diseases has much better coverage of funding information than the average level pre-2008 publications (around three times more), which enabled us to perform analysis on pre-2008 years. We acknowledge that the error rate of pre-2008 analysis would tend to be relatively larger, but we argue that the effect of this should not be substantial: we focus on correlations and not absolute numbers – unless there is a bias in the un-identified funded research pre-2008, this should only cause noise and not systematically affect our results. In addition, the research practices of scientists could also be an underlying factor that has made funded research less identifiable (Álvarez Bornstein et al., 2017), which highlights the need to better regulate research and funding practices.
3.2 Network Measures for ID
We defined a temporal network to model the evolution of the ID of infectious disease research, with the nodes being the bodies of infectious disease knowledge and links being their connection strength quantified by the cosine. To quantify ID, we studied the coherence and intermediation measures of the constructed networks. [^11] Note that every measure discussed in this section is computed given a time $t$, so we omitted index $t$ for simplicity.
We adopt a similar measure of coherence with Rafols and Meyer (2010), i.e., the mean linkage strength, due to its simplicity, although at a different level of aggregation. We used the mean node strength to represent the overall compactness of the knowledge system, i.e.,
$\overline{s}=\frac{1}{N}\sum_{i}s_{i}=\frac{1}{N}\sum_{i}\sum_{j\neq i}w_{ij}$ (1)
where $N$ represents the number of nodes in the network. Note that the node strength $s_{i}$ is a measure of how well an infectious disease integrates knowledge locally.
Following Leydesdorff (2007), we adopted betweenness centrality (BC) (Brandes, 2001) to track node-level ID from the intermediation perspective, i.e., one disease’s ability to bridge otherwise disjoint disease knowledge. BC of a node $i$ is computed as
$b_{i}=\sum_{j\neq k}\frac{g_{jk}(i)}{g_{jk}}$ (2)
where $g_{jk}$ is the number of shortest path between any two nodes and $g_{jk}(i)$ is the number of shortest path between two nodes going through node $i$.
Soós and Kampis (2012) proposed network diameter (or the maximal shortest path length) as a measure of both coherence and intermediation. This measure provides an upper bound of how quickly information flows from one place to another. We proposed a similar measure, the average shortest path length (ASPL) (Jahanshad et al., 2012), to capture the average shortest distance between infectious diseases. A low ASPL indicates that the disease network is compact. The shortest path length (Brandes, 2001) between node $i$ and $j$ is
$d(i,j)=\min\left(\frac{1}{w_{ih}}+\cdots+\frac{1}{w_{hj}}\right)$ (3)
with $h$ being intermediary nodes between node $i$ and $j$. ASPL is the average of $d(i,j)$ over all possible pairs of diseases $(i,j)$.
Beyond coherence and intermediation measures proposed in earlier literature, we were also interested in clustering patterns of the infectious disease knowledge network. For this purpose, following Blondel et al. (2008), we evaluated the networks modularity and used the Louvain method (Blondel et al., 2008) to determine modules of interconnected diseases. Modularity is computed as
$Q=\frac{1}{2m}\sum_{i,j}\left(w_{ij}-\frac{s_{i}s_{j}}{2m}\right)\delta(c_{i},c_{j})$ (4)
where $m$ is the sum of link weights, $s_{i}$ refers to the strength of node $i$, and $\frac{s_{i}s_{j}}{2m}$ is the expected link strength between $i$ and $j$ assuming a random distribution of connections which preserves the strength distribution across nodes. A high modularity $Q$ represents a well-defined community structure with many intra-community links and few links connecting separate communities while a low $Q$ indicates a weak community structure.
4 Results
4.1 Identification of temporal regimes in interdisciplinary research
Given two yearly slices of the correlation network at times $t_{1}$ and $t_{2}$, $w^{t_{1}}$ and $w^{t_{2}}$, we can quantify the distance between them by the Euclidean distance, i.e., $\sqrt{\sum_{i\neq j}(w_{ij}^{t_{2}}-w_{ij}^{t_{1}})^{2}}$. We then used bottom-up hierarchical clustering based on UPGMA (Unweighted Pair Group Method with Arithmetic mean) (Sokal and Michener, 1958) to partition the temporal evolution of publication patterns into distinct regimes. Further analysis of the obtained partitioning using the elbow method reveals three regimes for funded, unfunded and all research respectively (see Figure 1) which we identify with coherent periods of time, i.e. 1995-2007 (F1) [^12] We combined the single element clusters 1997 and 2004 with the bigger cluster to make F1 continuous as all years within F1 are already quite distant compared with that of F2 and F3. Years in F1 being more distant might be attributed to the larger noise in funded research as it is under-represented pre-2008 in the WoS, as discussed in Section 3.1. Visualisations with and without 1997 and 2004 were compared in Figure 1., 2008-2015 (F2), and 2016-2022 (F3) as regimes for funded research; 1995-2003 (U1), 2004-2015 (U2) and 2016-2022 (U3) for unfunded research; 1995-2003 (A1), 2004-2015 (A2), and 2016-2022 (A3) for all research [^13] Note that unfunded research has the same regimes as all research, so the following analysis will only be focusing on funded and unfunded regimes..
Figure 1: Bottom-up hierarchical clustering of yearly time slices for the funded, unfunded and aggregated networks. Hierarchical clustering using UPGMA was performed for all (A), funded (F), and unfunded (U) networks respectively. The highlighted identified significant clusters were based on the elbow method as in Figure 2 in the appendix.
Regimes for F and U were visualised in Figure 2. Each visualisation represents the average network in the corresponding regime. We observe regimes F1 and U1 tend to be relatively weakly connected compared with later regimes, indicating a weaker level of knowledge integration. This can also be seen from Table 1 where the average node strength $\overline{s}_{F}$ and $\overline{s}_{U}$ in F1 and U1 are 0.24 and 0.35, the lowest across all regimes. We note the existence of a well-integrated community of four curable sexually transmitted infections (STIs), i.e., chlamydia, gonorrhoea, syphilis and trichomoniasis in F1 and U1. The following links in F1 and U1 also demonstrate strong extents of integration: Hepatitis B (HBV) and Hepatitis C (HCV), dengue and yellow fever, varicella and herpes [^14] Varicella is linked with shingles as they are both caused by Varicella-Zoster Virus (VZV)., and diphtheria and tetanus (especially in U1).
Investigating the 2nd temporal regime, we observe that both F2 and U2 are more densely connected compared with the previous regimes, as can also be seen from $\overline{s}_{F}$ and $\overline{s}_{U}$ rising to 0.35 and 0.43, respectively, see Table 1. What stands out is that the knowledge integration among the three diseases of DTP is strongly reinforced.
Moving to F3 and U3, we observe that DTP gets integrated further, and the group of vector-borne (Aedes mosquitoes) diseases, i.e., yellow fever, zika and dengue, becomes relatively well integrated. Throughout the entire time period, the continuous integration of DTP knowledge is very likely caused by an increase in vaccination-related research. The sudden increase in the integration between zika, yellow fever, and dengue in the last regime is highly likely due to the increased attention on zika and its closely related neighbours on the knowledge network, as a reaction towards the emergence of zika as a public health emergency in 2016151515On 01 Feb 2016, WHO declared Zika and its complications constitutes a Public Health Emergency of International Concern. See the discussion for further details.. The integration of the four curable STIs (chlamydia, gonorrhoea, syphilis and trichomoniasis) and the group of hepatitis infections is found to be relatively stable, which might indicate an already well-established knowledge structure in those sub-fields.
Figure 2: Network visualisation of funded and unfunded temporal regimes. The link strength $w_{ij}$ shown in the visualisation of the regime is computed by the average of the link strength across the regime. Node size of disease $i$ represents the proportion of publications on $i$ during the regime. The network layout was produced using R based on the Fruchterman-Reingold method (Fruchterman and Reingold, 1991) applied on F1. The community structure was detected based on the Louvain method (Blondel et al., 2008).
We measure individual disease contributions to system change by characterising each disease’s relative impact in terms of relative link strength change $r_{s,within}$ versus the volatility $\sigma(diff[s_{i}(t)])$ in its time evolution of link strengths. More precisely, we defined $\sigma(diff[s_{i}(t)])$ to be the standard deviation of the change in node strength $s_{i}$ between consecutive years in the regime ($diff$ represents differencing of the time series), and $r_{s,within}$ to be the change of node strength from the beginning to the end of a regime normalised to the average level of change of all nodes, i.e., if a regime starts from year $t_{1}$ and ends in year $t_{2}$, $r_{s,within}=\frac{\overline{s_{i,t_{2}}}-\overline{s_{i,t_{1}}}}{\left|s_{t_{2}}-s_{t_{1}}\right|}$, where $\overline{s_{i,t_{2}}}$ represents the mean of $s_{i}$ across $t_{2}$, $t_{2}-1$, $t_{2}-2$, $\overline{s_{i,t_{1}}}$ represents the mean of $s_{i}$ across $t_{1}$, $t_{1}+1$, $t_{1}+2$, and $s_{t_{2}}$ and $s_{t_{1}}$ represents the average of $\overline{s_{i,t_{2}}}$ and $\overline{s_{i,t_{1}}}$ over all $i$ respectively. An increase in $r_{s,within}$ indicates the presence of relatively stronger local connections, thus being more integrated locally during the regime. In Figure 3, we plotted the relative link strength change $r_{s,within}$ on the y-axis versus the volatility $\sigma(diff[s_{i}(t)])$ on the x-axis. We observe that the changes in the different regimes seem to be driven by different diseases (Figure 3). Hepatitis A (HAV), chlamydia and dengue gain relatively more strength in F1 while yellow fever and three curable STIs (gonorrhoea, syphilis, and trichomoniasis) stand out in U1 (Figure 3). DTP emerges during F2 while varicella and pertussis stand out during U2. Throughout F3, there is a significant gain in the strength of coronavirus and zika, while for U3, coronavirus, tetanus and influenza stand out. All of the highlighted infectious diseases emerge with rather high volatility. The strong emergence of coronavirus-related bodies of knowledge in F3 and U3 is likely due to the huge attention paid to the COVID-19 global pandemic since 2019.
Figure 3: Analysis of contributions of individual diseases to system change within funded regimes (F1:1995-2007, F2:2008-2015, F3:2016-2022) and unfunded regimes (U1:1995-2003, U2:2004-2015, U3:2016-2022) in terms of relative change in node strength $r_{s,within}$ plotted against volatility $\sigma(diff[s_{i}(t)])$. The diseases with the highest $r_{s,within}$ were highlighted in each regime. Hepatitis A (HAV), chlamydia and dengue were highlighted in F1 while yellow fever, gonorrhoea, syphilis, and trichomoniasis in U1; DTP in F2 while varicella and pertussis in U2; coronavirus and zika in F3 while coronavirus, tetanus and influenza in U3.
4.2 Trends in compartmentalisation and integration of funded and unfunded research
In Table 1 we reported the average node strength $\overline{s}$, the modularity $Q$ and the average shortest path length $ASPL$ of the averaged networks for the regimes identified in Section 4.1. Comparing trends in $\overline{s}$, $Q$ and $ASPL$ for the different time regimes we make the following observations. First, regarding the average node strength $\overline{s}$, we find $\overline{s}_{F}$ increases from 0.24 to 0.35 to 0.44 and $\overline{s}_{U}$ from 0.35 to 0.43 to 0.54. This increase in the intensity of connection indicates that funded and unfunded research has become more interdisciplinary. Second, regarding the average shortest path length $ASPL$, we note a substantial decrease over time. The observed trend indicates a tendency towards higher interdisciplinarity in funded and unfunded research through stronger intermediation. Third, regarding the modularity $Q$, we find an opposite trend with $Q_{F}$ increasing and $Q_{U}$ decreasing, meaning that funded research becomes more compartmentalised while unfunded research becomes more globally integrated.
From the above, we infer that knowledge integration in funded research tends to reinforce community structure. This indicates that funded research tends to be more specialised and stays on the conservative side, i.e., it focuses on deepening already established relationships of diseases. Knowledge integration in unfunded research, on the other hand, tends to take place through weakening community structures, which indicates that unfunded research is relatively less conservative and focuses more on bridging distant diseases. These interpretations are further supported by results shown in Figure 3 and Figure 4, where we have compared relative increments in link strengths between time periods for the strongest (top 5%) and weaker (bottom 50%) links in the evolution of IDR in funded and unfunded research. We note, that the strongest links in funded research have consistently gained more in link strength compared to unfunded research (relative gains $4.96>3.74$ from 1995-2008 to 2009-2015 and even more so $4.13>2.67$ from 2009-2015 to 2016-2022). In contrast, for the majority of weaker links, links have gained more in strength for unfunded research than for funded research (relative gains $0.26>0.16$ from 1995-2008 to 2009-2015 and $0.22>0.15$ from 2009-2015 to 2016-2022). Both observations make it very clear that strength gains in funded research tend to be more aligned with already strong connections, whereas strength gains in unfunded research tend to be more exploratory, reinforcing weak connections.
Periods F1:1995-2007 F2:2008-2015 F3:2016-2022
$\overline{s}_{F}$ 0.24 0.35 0.44
$Q_{F}$ 0.46 0.47 0.49
$ASPL_{F}$ 127 98 89
Periods U1:1995-2003 U2:2004-2015 U3:2016-2022
$\overline{s}_{U}$ 0.35 0.43 0.54
$Q_{U}$ 0.45 0.43 0.40
$ASPL_{U}$ 92 77 64
Table 1: Average node strength $\overline{s}$, modularity $Q$ and average shortest path length $ASPL$ for funded (F) and unfunded (U) regimes.
4.3 Analysing research funding into Interdisciplinary Research
In Section 4.2, observing that in contrast to unfunded research funded research tended to become more compartmentalised, we noted different trends in the organisation of funded and unfunded research over time. Here, we are interested in a more complete understanding of how research investment has driven the evolution of interdisciplinarity. For this purpose, one could see unfunded research as representing general scientific interest and compare this to research driven by funding allocation.
To operationalise this comparison, for each temporal regime, we measured average correlations between pairs of diseases in funded and unfunded research and plotted them against each other in Figure 4. Note that in Figure 4 we show results based on partitions for the temporal regimes in both U and F (top and bottom rows), which show essentially the same trends. The 45-degree line in Figure 4 represents a situation where research corresponding to a pair of diseases is as well-funded as represented in general scientific interest, i.e. in unfunded research. Inspecting the figure, we first observe that most of the pairs stay below the 45-degree line, indicating that the level of interdisciplinarity in funded research is generally lower than in unfunded research. In other words, IDR in infectious disease research tends to be underfunded. However, this changes over time as the slope of the fitted regression line keeps increasing and approaches one, getting closer to where funding allocation matches overall scientific attention. Simultaneously, also the R-squared value increases over time, indicating an increasingly closer alignment between research investment and scientific interest regarding the importance of interdisciplinary infectious disease areas.
Further to the above, Figure 4 also allows to analyse how well a pair of diseases is funded relative to all other pairs. Visually, such relative over- or underfunding is indicated by whether the corresponding datapoint is above or below the fitted regression line. For instance, in Figure 4, we highlight seven pairs of diseases that received the most scientific attention in the last temporal regime. Looking at trends over time, we note that in F1 and U1, the research area chlamydia-gonorrhoea stays above the fitted regression line indicating a relatively well-funded status, whereas the DTP-related areas stay slightly below the line indicating they are relatively underfunded. These observations remain true for F2 and U2, except that HCV-HBV becomes slightly underfunded and DTP-related areas move slightly closer to the fitted line. In F3 or U3, dengue-zika becomes very well-funded relative to other areas whereas HCV-HBV becomes relatively underfunded.
Figure 4: The evolution of the level of knowledge integration in funded research compared with unfunded research for (A) U1:1995-2003, U2:2004-2015 and U3:2016-2022; and (B) F1:1995-2007, F2:2008-2015 and F3:2016-2022. For each plot, the y-axis represents the average link strength of a pair of infectious diseases in funded research and the x-axis in unfunded research. Each error bar represents the standard error of the link strength of a pair within the regime. The red line represents the 45-degree line: any link lying on the line represents the same level of ID in funded and unfunded research, above (below) indicates more (less) funding is allocated to the pair than should be. A regression line is fitted to the points and the slope of the fitted regression line, the standard error of the slope, and the R-squared value of the fitted regression line were reported in the top left corner. The shaded area around the fitted line represents one standard error of the slope. Yellow labels represent the top seven influential links that appeared in F3 or U3, and their corresponding positions were shown in F2, F1, U2, and U1 plots.
4.4 Comparing the roles of diseases in knowledge integration
To identify the diseases with the strongest global and local impact, we ranked the diseases by strength and betweenness. Table 2 reports the top three ranked strengths $s_{i}$ and betweenness centrality $b_{i}$ of the infectious diseases for F and U regimes. $s_{i}$ represents the knowledge integration at a local level and $b_{i}$ at a global level, thus comparing these two allows us to compare the local and global impact of the top-ranked diseases. Comparing the different time regimes we make the following observations. Regarding the betweennees centrality $b_{i}$, we find HIV has the highest $b_{i}$ in all regimes for both F and U, followed by other prominent diseases like tuberculosis, HBV, or malaria. This indicates that prominent infectious diseases like HIV stay in the centre of the network and bridge distant bodies of knowledge. Regarding $s_{i}$, we find DTP are the top three in $s_{i}$ in F2, F3, U2 and U3 with a remarkable increase in magnitude through time, but none of DTP appears in the top three $b_{i}$ of F or U regimes. This indicates that DTP stay on the periphery of the network and only enhance established connections between bodies of knowledge.
Regimes F1:1995-2007 F2:2008-2015 F3:2016-2022
Top Three $s_{i}$Centerstack HIV(0.59)
Gono(0.51)
[…]
Diph(0.86)
[…]
Diph(1.17)
Pert(1.03)
Top Three $b_{i}$Centerstack HIV(0.44)
TB(0.24)
[…]
Malaria (0.36)
[…]
Malaria(0.24)
Pneum(0.15)
Regimes U1:1995-2003 U2:2004-2015 U3:2016-2022
Top Three $s_{i}$CenterstackDiph(0.86)
Teta(0.83)
[…]
Teta(1.19)
[…]
Teta(1.41)
Pert(1.32)
Top Three $b_{i}$CenterstackHIV(0.30)
HBV(0.29)
[…]
Malaria(0.22)
[…]
Malaria(0.17)
Dengue(0.16)
Table 2: Top three ranked diseases, ranked by node strength $s_{i}$ and betweenness centrality $b_{i}$ for the funded (F) and unfunded (U) regimes.
We investigated the relative change in the diseases’ betweenness $r_{b,within}$ with respect to their volatility $\sigma(diff[b_{i}(t)])$ in Figure 5. $r_{b,within}$ and $\sigma(diff[b_{i}(t)])$ were calculated in analogy to $r_{s,within}$ and $\sigma(diff[s_{i}(t)])$ in Section 4.1. Note that $r_{b,within}$ for each infectious indicates a change in its ability to act as a bridge between other diseases relative to the average change in betweenness of all infectious diseases within a regime. From Figure 5, we note that HCV, HAV and HIV have gained notable betweenness during F1, while dengue and HAV have gained the most betweenness during U1. Malaria has gained strongly in betweenness in F2 with high volatility, whereas during U2 pertussis and varicella have gained betweenness with moderate volatility. Throughout F3, coronavirus has gained significant betweenness with moderate volatility, while during U3 it is ebola, pertussis and measles that have gained notable betweenness. The significant gain of coronavirus is likely due to the COVID-19 pandemic, but interestingly this pattern is not found in unfunded research.
Figure 5: Analysis of contributions of individual diseases to system change within F regimes (F1:1995-2007, F2:2008-2015, F3:2016-2022) and U regimes (U1:1995-2003, U2:2004-2015, U3:2016-2022) in terms of relative change in betweenness $r_{b,within}$ versus volatility $\sigma(diff[b_{i}(t)])$. The diseases with the highest $r_{b,within}$ were highlighted in each regime. HCV, HAV and HIV were highlighted in F1 while dengue and HAV in U1; malaria in F2 while varicella and pertussis in U2; coronavirus in F3 while measles, ebola and pertussis in U3.
4.5 Quantifying the Role of Coronavirus Research in IDR
To further explore the role of the coronavirus pandemic in infectious disease research, in Figure 6(a) we reported the changing ranking of coronavirus research in terms of the number of publications, node strength and betweenness over time. We also marked two important coronavirus-related public health events in the figure, i.e., the 2002-2004 SARS outbreak (yellow regions) and the 2019-2022 COVID-19 pandemic (red regions).
For both the SARS and COVID-19, we observe a significant rise in the publication ranking of both F and U. During SARS the publication ranking of coronavirus broke into the top 15 for F and top 10 for U, while during COVID-19 it attained the top rank for both F and U. Unsurprisingly, we see that both events led to an increase in scientific attention, with COVID-19 to the greatest extent. This is consistent with the result in Figure 2 where the proportion of coronavirus-related research (represented by the node size) outweighs HIV and becomes the highest in F3 and U3. However, in terms of the ranking of node strength, despite experiencing moderate ranking gains during SARS and COVID-19, coronavirus stayed out of the top 15 for both F and U. This might indicate that the outbreaks caused some local knowledge integration around coronavirus but not at a significant extent at the system level.
Moreover, in terms of the betweenness ranking, U exhibited only minor fluctuations around the 20th to 25th during both events, while F showed a minor drop around the 20th rank during SARS but a considerable jump from the 30th to a top 10 position during COVID-19. Such a jump might indicate coronavirus has become increasingly important by moving more into the centre of the infectious disease network and starting to bridge distant knowledge. However, even though coronavirus has had a very important role in terms of the number of publications, its systemic impact on the interdisciplinarity of infectious disease research has been relatively small to date. This might be due to the fact that coronavirus is conceptually not so strongly related to other disease areas. Another potential reason is there perhaps exists a temporal delay before systemic impact is observed.
To explore the potential for delays in the systemic impact of COVID-19 on coronavirus research, we compared to another infectious disease with a sudden increase in prominence for which a longer timeframe of observations is available. This is provided by the zika virus which strongly gained in attention during the global outbreak between 2015 and 2016. For both F and U, strength seemed to immediately follow publication ranking during the outbreak 2015-2016 (Figure 6(b)), which is different to COVID-19. The betweenness ranking in F lagged a bit behind but then became quite volatile, while in U there seemed no delay. We conclude there is no sufficient evidence for a delay in the systemic impact of the zika outbreak. Funded research of zika has demonstrated extraordinary local integration (perhaps with other vector-borne diseases) with strength ranking being the best performer among all three rankings (peaked at the top 5 and sustained at that level ever since). This result on funded research on zika shows consistency with results in Section 4.1, Section 4.2, and Section 4.3.
Figure 6: Temporal change in the ranking of the annual number of publications, strength and betweenness of (a) coronavirus-related publications (b) zika-related publications in the funded and unfunded disease network 1995-2022. The yellow area in (a) indicates the SARS outbreak 2002-2004, and the red area in (a) indicates the COVID-19 outbreak 2019-2022. The red area in (b) indicates the zika outbreak 2015-2016.
5 Discussion
We summarise the results and discuss the practical implications of this study in Section 5.1, then discuss real-world factors behind the results, limitations and future work in Section 5.2.
5.1 Summary of results and practical implications
We have investigated the evolution of interdisciplinarity in funded and unfunded research on infectious diseases over the period 1995-2022. Constructing correlation networks of research output relating to pairs of diseases, we identified three regimes for funded and unfunded research respectively where each regime is a coherent period of time characterised by a particular knowledge structure. Based on the regimes, we found that both the funded and unfunded research had an increase in the extent of knowledge integration through time in terms of coherence and intermediation. However, while increases in interdisciplinarity in funded research took place through compartmentalisation, increases in unfunded research have typically occurred through global integration. Besides, we also found that IDR on infectious disease is underfunded in general but also note that this effect has decreased through time. Our analysis further allows to identify individual well-funded and underfunded interdisciplinary areas.
Investigating the role of individual diseases in these trends, we found IDR on prominent diseases like HIV, malaria and tuberculosis has strong bridging effects, while IDR on diphtheria, tetanus, and pertussis has strong local enhancement. Lastly, we found that coronavirus has attracted the most publications in infectious disease research since the emergence of COVID-19. In spite of this, however, the systemic impact of coronavirus research to date on infectious disease knowledge integration has been relatively small.
The results of this study contribute to the understanding of the roles played by the research on individual infectious disease in interdisciplinary knowledge generation, and the relationship between global public health emergencies and interdisciplinary research efforts. These could provide valuable insights for future health priority setting and resource allocation.
The proposed analytical framework of funded and unfunded IDR can be applied to any field of science. It could provide valuable insights to researchers, funders and policymakers on the evolution of IDR and the over/under-funded interdisciplinary areas, supporting funding decisions to facilitate IDR.
5.2 Real-world factors, limitations, and future work
The observed compartmentalisation in funded research output in terms of ID is consistent with the lack of appreciation towards IDR during research evaluations (Bromham et al., 2016; Boudreau et al., 2016) and the observed conservativeness of research funding (Packalen and Bhattacharya, 2020). Funding sources, especially those that routinely issue topic-focused calls, seem to have the greatest influence on most science and health research which are resource intensive (Thelwall et al., 2023; Whitley et al., 2018); however, the degree of such influence depends on individual funders’ requirements (Serrano Velarde, 2018; Luukkonen and Thomas, 2016; Thelwall et al., 2023). The observation that unfunded research is more widely exploratory in terms of ID is consistent with the fact that unfunded researchers have a higher degree of freedom in setting their own goals (Edwards, 2022).
The rising consistency between funded and unfunded research might suggest a better allocation of funding in the sense that interdisciplinary research areas with high scientific interest have been addressed by research funding. However, this would need further in-depth examination as we did not take into account the complex interplay between funded and unfunded research in our study. We want to also note that the focus of our work is on topic-level ID instead of ID based on conventional classification systems. The granularity of the choice of classifications is likely to have an impact on the resulting ID (Rousseau et al., 2019). We leave in-depth examinations of such impact for future work.
We distinguished (the research on) infectious diseases with two different patterns of knowledge integration: diseases that tend to integrate locally (e.g. DTP) and diseases that tend to integrate globally (e.g. HIV, tuberculosis, and malaria). There is no simple answer to what drives these integrations as the collection of knowledge on infectious diseases spans multiple domains including pathogenesis, diagnosis, cause, treatment, prognosis, spread, prevention, social impact and policy. However, there may be fundamentally two types of driving factors: intentional or unintentional (Wagner et al., 2011; Glänzel and Debackere, 2022). Policymakers, funders and researchers have been targeting specific groups of diseases to meet their respective priorities and goals: these are intentional efforts, and they tend to focus on more established groups of diseases that have been associated with great disease burdens at that time. For instance, the WHO has created global health programmes focusing on Hepatitis B and C [^16] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/hepatitis/overview (Date last accessed: 2024.9.4), STIs [^17] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/stis/overview (Date last accessed: 2024.9.4), and vector-borne infections [^18] https://www.who.int/publications/i/item/9789241512978 (Date last accessed: 2024.9.4). Therefore, it might be that emergence of the diseases with strong local integration is the result of such intentional effort. For example, DTP, the diseases found with the largest extent of local integration, have been a part of the WHO Expanded Programme on Immunization [^19] https://www.who.int/teams/immunization-vaccines-and-biologicals/essential-programme-on-immunization (Date last accessed: 2024.9.4), and a focus of Gavi, the Vaccine Alliance [^20] https://www.gavi.org/news/media-room/gavi-helps-dtp3-coverage-rise-after-stagnation (Date last accessed: 2024.9.4), and there has been ongoing development of a variety of DTP-related combination vaccines [^21] https://www.who.int/teams/health-product-policy-and-standards/standards-and-specifications/norms-and-standards/vaccine-standardization/dt-based-combined-vaccines (Date last accessed: 2024.9.4).
Unintentional factors like discoveries on biological associations, patterns of comorbidity, or a technology or knowledge spill-over across diseases could all play a role in advancing knowledge integration. HIV, tuberculosis, and malaria being the main drivers of global integration, account for 52.1% of total infectious disease funding from G20 countries in 2000-2017 while HIV alone accounts for 40.1% of total funding with 42.1 billion US dollar (Head et al., 2020). The substantial resources specifically devoted to HIV research have led to collateral benefits to other disease areas (Schwetz and Fauci, 2019). Some examples include advancing antiviral drug development (on hepatitis C), improving vaccine research techniques (on ebola, zika, and influenza), enhancing understanding of immunology (on the role of $CD4+$ $T$ cells in fighting other infectious diseases and certain cancers) and advancing structural biology (on structure-based vaccine design that can be applied to other pathogens) (Schwetz and Fauci, 2019). Such unintentional but broad spin-offs might have enabled research on prominent diseases like HIV to bridge a wide range of bodies of knowledge and drive knowledge integration in the field of infectious disease research. Further disentanglement of different mechanisms behind integration might require identifying research fields and types of science of the publications on infectious diseases, and we leave this task to future research.
We also found that the number of publications on coronavirus has skyrocketed since COVID-19 emerged. The generated scientific knowledge of coronavirus research has been informing public health responses, treatments, and vaccine development (Micah et al., 2023). However, it has been argued that this surge reflects opportunism by both researchers and journals (Clark, 2023). For researchers, there has been a “covidisation” of research to remain relevant and secure funding; for journals, there has been a loose “gate-keeping” followed by fraudulent and poor quality research but eventually higher impact factors due to bulk citations (Clark, 2023; Glasziou et al., 2020). Despite the surge in publications, the systemic impact of coronavirus on IDR was found to be fairly small. We suspected there might be a temporal delay for the systemic impact to catch up, but after validating the idea on the zika outbreak we found no evidence for such a delay. We encourage future research to further investigate this in depth.
CRediT authorship contribution statement
Anbang Du: Conceptualization, Software, Data curation, Resources, Formal Analysis, Methodology, Visualization, Writing – original draft, Writing – review & editing
Michael Head: Supervision, Data curation, Resources, Writing – review & editing
Markus Brede: Conceptualization, Supervision, Methodology, Writing – review & editing
Conflict of interest
There is no conflict of interest for this manuscript.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Data availability
Data and code are available on request.
[…]
{xltabular}
[…]
Coronavirus ”COVID” OR ”COVID-19” OR ”Coronavirus” OR ”Corona virus” OR ”2019-nCoV” OR ”SARS-CoV” OR ”MERS-CoV” OR ”Severe Acute Respiratory Syndrome” OR ”Middle East Respiratory Syndrome” 284084 120371
HIV ”HIV” OR (”AIDS” AND (immun* OR patient* OR epidem* OR pandemic*)) OR ”Human immunodeficiency virus” OR ”Acquired Immune Deficiency syndrome” OR ”acquired immunodeficiency syndrome” 277242 143086
Pneumonia ”Pneumonia” OR ”pneumonias” OR ((lower respiratory tract infection*) OR (severe respiratory tract infection*)) 106260 40569
Tuberculosis ”Tuberculosis” 102360 48078
Influenza (”flu” AND (pandemic* OR vaccin* OR shot* OR season*)) OR ”influenza” 74697 44141
Hepatitis C ”Hepatitis C” OR (”hcv” AND (infect* OR virus* OR patient* OR hepatitis OR liver)) 6829528374
Malaria ”Malaria” OR ”Malarial” OR Plasmodium infect* 65608 39227
Salmonella ”Salmonella” 61690 31563
Hepatitis B ”Hepatitis B” OR (”hbv” AND (infect* OR virus* OR patient* OR hepatitis OR liver)) 57374 26185
Herpes (”HSV” AND (infect* OR vaccin* OR 1 OR 2 OR virus*)) OR ”Herpes” OR ”Shingles” 37844 17525
Urinary Tract Infection ”Urinary Tract Infect*” OR (”UTI” AND (E. coli OR antibiotic OR chlamydia OR Patient*)) 30919 10037
Meningitis”Meningitis”274758692
Dengue ”Dengue” 2263913916
Chlamydia ”Chlamydia” OR ”Chlamydiae” OR ”Chlamydial” 17801 8137
Leishmaniasis ”leishmaniasis” 174219651
Pertussis ”Pertussis” OR ”whooping cough” 15864 6943
Measles ”Measles” 11464 4622
Tetanus ”Tetanus” 108574462
Chagas ”chagas” OR ”American trypanosomiasis” 10684 6126
Syphilis ”syphilis” 10313 3810
Varicella ”Varicella” OR ”Chickenpox”90493032
Schistosomiasis ”Schistosomiasis” 88334685
Zika ”Zika” 8182 6129
Rabies ”Rabies” 7935 3540
Ebola ”Ebola” OR ”Ebolavirus” 75594565
Hepatitis A ”Hepatitis A” OR (”hav” AND (infect* OR virus* OR patient* OR hepatitis OR liver))7458 2367
Diphtheria ”Diphtheria” 72993536
Leprosy ”Leprosy”69632262
Hepatitis E ”Hepatitis E” OR (”hev” AND (infect* OR virus* OR patient* OR hepatitis OR liver))50802747
Gonorrhoea”N gonorrhoeae” infect* OR ”Neisseria gonorrhoeae” Infect* OR ”Gonorrhoea”49322543
Yellow fever”yellow fever”42452464
filariasis (filaria* AND (lymph* OR Elephantia* OR BANCROFTI* OR MALAYI OR Brugia*))37231972
trypanosomiasis ”Sleeping Sickness” OR ”African trypanosomiasis” OR ”Trypanosoma brucei gambiense” OR ”Trypanosoma brucei rhodesiense”35712301
Scabies ”Scabies”2126615
Onchocerciasis ”Onchocerciasis”1730802
Trichomoniasis ”Trichomoniasis”1569770
Search terms, total and funded number of records for each disease from 1995 to 2022. The data extraction was performed on 2024.03.29.
{xltabular}
[…]
Corona Coronavirus
HIV HIV
Pneum Pneumonia
TB Tuberculosis
Influenza Influenza
HCV Hepatitis C
Malaria Malaria
Salm Salmonella
HBV Hepatitis B
Herpes Herpes
UTI Urinary Tract Infection
Mening Meningitis
Dengue Dengue
Chlamy Chlamydia
Leishma Leishmaniasis
Pert Pertussis
Measles Measles
Teta Tetanus
Chagas Chagas
Syphilis Syphilis
Vari Varicella
Schisto Schistosomiasis
Zika Zika
Rabies Rabies
Ebola Ebola
HAV Hepatitis A
Diph Diphtheria
Leprosy Leprosy
HEV Hepatitis E
Gono Gonorrhoea
YF Yellow Fever
Fila Filariasis
Trypano Trypanosomiasis
Scabies Scabies
Onchocer Onchocerciasis
Trichomo Trichomoniasis
Disease names and abbreviations used in this study.
Year WOS F36 FAll WOS F WOS All 36 F 36
1995 0.06 0.18 798461 43966 18818 3372
1996 0.05 0.17 886477 47519 22566 3801
1997 0.05 0.17 898426 47278 22952 3853
1998 0.05 0.17 923073 47735 23595 4097
1999 0.05 0.17 897754 48810 24177 4161
2000 0.05 0.17 939156 47548 24608 4063
2001 0.05 0.16 933384 48926 24073 3922
2002 0.05 0.16 951628 50729 24596 3910
2003 0.05 0.16 1008291 54242 25677 4156
2004 0.06 0.17 1050680 59610 27334 4727
2005 0.06 0.19 1179694 75764 29187 5622
2006 0.11 0.21 1251501 140190 31235 6520
2007 0.12 0.21 1371632 158111 33888 7168
2008 0.22 0.34 1527949 328555 36557 12463
2009 0.39 0.53 1638275 635089 39573 21012
2010 0.43 0.58 1688059 730003 43165 25125
2011 0.46 0.61 1796994 820457 46694 28267
2012 0.47 0.62 1902726 893601 48418 30011
2013 0.48 0.64 2000048 960986 50688 32378
2014 0.48 0.64 2108895 1007809 51995 33419
2015 0.48 0.65 2191395 1059818 53713 35027
2016 0.51 0.65 2295920 1165592 55004 35863
2017 0.57 0.67 2381988 1354334 56551 37834
2018 0.58 0.67 2442581 1416402 56792 38025
2019 0.59 0.66 2644412 1564103 59708 39574
2020 0.60 0.52 2795066 1672916 110972 57726
2021 0.62 0.51 3023128 1871583 181063 91837
2022 0.62 0.51 3078441 1911288 188602 95997
Table 1: The number and proportion of funded publications through time in the WOS Core Collection. All WOS: the total number of WoS publications (filtered by SQ1). F WOS: all funded WoS publications (filtered by SQ1 and SQ2). WOS F%: the proportion of funded WoS publications. All 36: the total number of WoS publications on the 36 diseases (filtered by SQ1 and all 36 diseases’ search terms joined by ’OR’ in the Topic field). F 36: the total number of WoS publications on the 36 diseases (filtered by SQ1 and SQ2 and all 36 diseases’ search terms joined by ’OR’ in the Topic field). 36 F%: the proportion of funded WoS publications on the 36 diseases. Data extraction was done on 10/04/2024. The drop in 36Fstarting the year 2020 was due to the mass increase in coronavirus publication coupled with a low rate of identified funded research as shown in Table 2.
Year Corona Corona F Corona 36% Corona F36 F%
1995 116 32 0.01 0.28 0.18
1996 96 24 0.00 0.25 0.17
1997 112 27 0.00 0.24 0.17
1998 138 28 0.01 0.20 0.17
1999 101 23 0.00 0.23 0.17
2000 95 20 0.00 0.21 0.17
2001 122 29 0.01 0.24 0.16
2002 82 15 0.00 0.18 0.16
2003 325 33 0.01 0.10 0.16
2004 775 59 0.03 0.08 0.17
2005 711 91 0.02 0.13 0.19
2006 615 122 0.02 0.20 0.21
2007 454 101 0.01 0.22 0.21
2008 447 166 0.01 0.37 0.34
2009 393 259 0.01 0.66 0.53
2010 372 247 0.01 0.66 0.58
2011 310 225 0.01 0.73 0.61
2012 308 231 0.01 0.75 0.62
2013 380 288 0.01 0.76 0.64
2014 474 349 0.01 0.74 0.64
2015 497 330 0.01 0.66 0.65
2016 541 366 0.01 0.68 0.65
2017 553 410 0.01 0.74 0.67
2018 507 391 0.01 0.77 0.67
2019 583 441 0.01 0.76 0.66
2020 48666 16710 0.44 0.34 0.52
2021 116651 49424 0.64 0.42 0.51
2022 127050 57309 0.67 0.45 0.51
Table 2: The number and proportion of funded Coronavirus publications through time in the WoS. Corona: total number of Coronavirus-related publications (filtered by SQ1). Corona F: total number of Coronavirus-related funded publications (filtered by SQ1 and SQ2). Corona 36%: proportion of Coronavirus-related publications within all WoS publications on the 36 diseases. Corona F%: proportion of funded Coronavirus-related publications out of all Coronavirus-related publications. 36 F%: the proportion of funded WoS publications on the 36 diseases. Data extraction was done on 16/04/2024.
Figure 1: Comparison of the average funded network 1995-2007 including or not including(*) the years 1997 and 2004. There is no major difference in the network structure after including these two years, except that certain links get slightly weakened due to the noise introduced.
Figure 2: Elbow method of identifying the optimal number of clusters. The within-cluster sum of squares (WSS) for different cluster numbers is measured, where the WSS of a cluster is its sum of squares to the centroid, i.e., WSS $=\sum_{i=1}^{k}\sum_{x\in C_{i}}|x-c_{i}|^{2}$, with $C_{i}$ the set of years belonging to cluster $i$’s and $c_{i}$ the clusters centroid. The funded elbow plot is created by removing the years 2004 and 1997. We observe that choosing three clusters would be a suitable choice for funded, unfunded, and all research.
Figure 3: Comparison of relative increments of the top 5% strongest links. Bar height represents the average link strength of the corresponding disease pair in the first period, and the solid line represents the pair’s increment in strength to the second period, relative to the average change in link strength. The dashed lines give average relative increments for all pairs included in the figure for funded (blue) and unfunded (orange) research. (a) Change from F1 (1995-2008) to F2 (2009-2015) with average relative increments of 4.95 for funded and 3.74 for unfunded research. (b) Change from F2 (2009-2015) to F3 (2016-2022), with average relative increments of 4.13 for funded and 2.67 for unfunded research.
Figure 4: Comparison of relative increments of the bottom 50% weakest (non-zero) links. Bar height represents the average link strength of the corresponding disease pair in the first period, and the solid line represents the pair’s increment in strength to the second period, relative to the average change in link strength. The dashed lines give average relative increments for all pairs included in the figure for funded (blue) and unfunded (orange) research. (a) Change from F1 (1995-2008) to F2 (2009-2015), with average relative increments of 0.16 for funded and 0.26 for unfunded research. (b) Change from F2 (2009-2015) to F3 (2016-2022), with average relative increments of 0.15 for funded and 0.22 for unfunded research.